OpenStudy (anonymous):

Which value of x gives the solution of 1/2 (6 + 4x) = x + 2(x + 4)?

OpenStudy (anonymous):

??????

OpenStudy (anonymous):

distribute first \[3+2x=x+2x+8\]

OpenStudy (anonymous):

then combine like terms \[3+2x=3x+8\] then finish in two steps

OpenStudy (anonymous):

so 5x=11x?